# Attacking at non-harmonic frequencies in screaming-channel attacks

### v1

@@ 7 Discussion and conclusion
This study also demonstrates how this phenomenon can make attacks feasible in cases where all exploitable harmonics are polluted by interfering signals, as the case in more realistic, real-life scenarios. The studied phenomenon can also reduce the number of traces needed for the attack. Compared with the performance of the attack at the best harmonic, using the best non-harmonic enables to reduce by $60\%$ the number of traces needed to get a key rank under $2^{35}$.
In future works, it could be interesting to detect the best frequencies first and then focus the efforts only on them. For example, by building better profiles: in our work, in order to build profiles at a large number ($150$) of frequencies, these profiles were built with only $150$K traces ($15000\times 10$), which is relatively small compared with previous works ($1$ to $5$M traces per profile). One can also extend the range of attacks to a distance where no harmonic gives a reasonable key rank (for example, superior to $2^{39}$) with a given maximum number of traces and observe how many meters a non-harmonic attack is capable of gaining.

### v2

@@ 7 Discussion and conclusion
This study also demonstrates how this phenomenon can make attacks feasible in cases where all exploitable harmonics are polluted by interfering signals, as is the case in more realistic, real-life scenarios. The studied phenomenon can also reduce the number of traces needed for the attack. Compared with the performance of the attack at the best harmonic, using the best non-harmonic enables to reduce by $60\%$ the number of traces needed to get a key rank under $2^{35}$.
In future works, it could be interesting to detect the best frequencies first and then focus the efforts only on them. For example, by building better profiles: in our work, in order to build profiles at a large number ($150$) of frequencies, these profiles were built with only $150$K traces ($15000\times 10$), which is relatively small compared with previous works ($1$ to $5$M traces per profile). One can also extend the range of attacks to a distance where no harmonic gives a reasonable key rank (for example, superior to $2^{39}$) with a given maximum number of traces and observe how many meters a non-harmonic attack can gain.
Acknowledgment
We want to acknowledge the reviewers of the current and previous versions of this paper, as well as Dr. Maria Méndez Real and Dr. Dennis Gnad for their constructive feedback.